I cannot do this im stuck...
Cashews sell for $8 a pound,and peanuts for $3 a pound. A store wants to make 20 pounds of mixed nuts which they will sell for $6 a pound. How many pounds of cashews and how many pounds of peanuts should they use?

Answer by mananth(16946)

let the quantity of cashews added be x
the quantity of peanuts to be added will be 20-x
the value of each added = the value of total
3*(20-x)+8x=20*6
60-3x+8x=120
5x=60
x=12 lbs cashews
y=8 lbs peanuts